AO Commissioning SX

Sun Oct 23 00:42:49.625 2016 aos.AdsecSet.failed - AdsecSet [18] - SetFlatAo: FATAL - IDL An in-use actuator used is not working anymore

Sun Oct 23 00:43:15.207 2016 aos.AdsecSet.started - AdsecSet [19]

Sun Oct 23 00:43:15.225 2016 aos.AdsecSet.failed - AdsecSet [19] - SetFlatAo: WARNING - adsecarb:Illegal command for state Failure (Set Flat AO)

It set fine on the second try.

fastdiagn.L |ERR| 16142|2016-10-23 00:52:55.296806| FUNCTEMERGENCYST > emergencyReact = 1 [Funct.cpp:95]
fastdiagn.L |INF| 16146|2016-10-23 00:52:55.323837| FUNCTEMERGENCYST > [FunctEmergencyStop] try to send Alert to AdSec Arbitrator [Funct.cpp:124]
fastdiagn.L |INF| 16147|2016-10-23 00:52:55.323940| FUNCTEMERGENCYST > [FunctEmergencyStop] Alert to AdSec Arbitrator succesfully sent! [Funct.cpp:134]

Unknown rip adsec was in a funny state.

Sun Oct 23 01:19:12.916 2016 aos.OffsetXY{FLAO}.failed - OffsetXY {FLAO}[37] - [AOException] RETRY: Offset outside FoV limits. Target xyz is -437.41 -51.54 54.59 (-5002) VALUE_OUT_OF_RANGE_ERROR

Only the large offset.

wfsarb.L |INF| 311|2016-10-23 01:26:40.564789| WFS > Python command: WfsArbScripts.offsetXY(self, sensor = 'FLAOWFS', instr = 'LUCIFER', offsetx = 8.37766, offsety = 0.013775, brake = True)
wfsarb.L |INF| 312|2016-10-23 01:26:47.334301| WFS > Command returns success .
wfsarb.L |INF| 313|2016-10-23 01:26:50.826562| WFS > Executing resumeLoop...
wfsarb.L |INF| 314|2016-10-23 01:26:50.826604| WFS > Python command: WfsArbScripts.resumeLoop(self, sensor = 'FLAOWFS', instr = 'LUCIFER')wfsarb.L |WAR| 315|2016-10-23 01:26:50.873459| WFS > Command returns retryable error : RETRY: Magnitude calculation cannot take place with negative counts (-5002) VALUE_OUT_OF_RANGE_ERROR

I've never seen that one before.

It might be that we missed the star, but I don't see how.

01:38 Variable seeing, using "Bad Seeing Table".

01:41 Offsets working. NCPA and GOPT working as they should be.

02:44 Everything is running smoothly, but there is a lot of vibration.

03:57 Finished the entire script!!! (P-Cygni)

03:58 Starting BIN 2 script.

04:32 It was running in LN gains for some reason, I re did the gains manually. It was in the low gains for maybe two images. I saw the telescope doing offsets like normal, but the low order gains were applied.

04:41 Still a lot of vibrations

04:59 Finished

-- %USERSIG{GregoryTaylor - 2016-10-23}%
Topic revision: r3 - 23 Oct 2016, GregoryTaylor
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback